OpenStudy (anonymous):

Peter earns $32,000 per year plus a 2.5% commission on his jewelry sales. Find Peter's total salary for the year when his sales are valued at $420,000.

OpenStudy (bahrom7893):

420,000 - 100% x - 2.5% x = 2.5*420,000/100 = 2.5*4,200 = 10500

OpenStudy (bahrom7893):

Total: $32,000+$10,500 = $42,500

OpenStudy (anonymous):

thank youu:)